completey sur ounds a colection of charges. Pind the electric nux te th its sign) through the arte rthe tolection conssts ara. ing s4S, , i Cth 4.80 x 10 C charge, and (c) both of the charges in (a) and (b) (a) Numbe [935 Units |N-mA2/C 5.42E5 4.5
Answer #1

(c)for both the charges in a and b

Q = Qa + Qb

Q = (8.3 - 4.8) x 10^-6 = 3.5 x 10^-6

phi = 3.5 x 10^-6/(8.85 x 10^-12) = 3.95 x 10^-5 N m^2/C

Hence, phi = 3.95 x 10^-5 N m^2/C